OpenStudy (anonymous):

Solve the equation for x: 7x – 4y = 6 + 2x.

OpenStudy (anonymous):

you need x in terms of y ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok subtract 2x from both of equations 7x-2x-4y=6+2x-2x 5x-4y=6 ok now add 4y on both sides

OpenStudy (anonymous):

so whats the final answer

OpenStudy (anonymous):

let me know what u got after adding 4y on both sides

OpenStudy (anonymous):

the answer is supposed to be a fraction an im not getting a fraction

OpenStudy (anonymous):

adding 4y on both sides 5x-4y+4y=6+4y 5x=6+4y

OpenStudy (anonymous):

divide both sides by 5 and tell me what you got :?

OpenStudy (anonymous):

im still not getting a fraction

OpenStudy (anonymous):

yes you are ! divide both sides by 5 \[\frac{5x}{5}=\frac{6+4y}{5}\] or \[x=\frac{6+4y}{5}\]

OpenStudy (anonymous):

got it ?